### v1

# The geometry of preperiodic points in families of maps on $\mathbb{P}^{N}$

Abstract.
We study the dynamics of algebraic families of maps on $\mathbb{P}^{N}$, over the field $\mathbb{C}$ of complex numbers, and the geometry of their preperiodic points. The goal of this note is to formulate a conjectural characterization of the subvarieties of $S\times\mathbb{P}^{N}$ containing a Zariski-dense set of preperiodic points, where the parameter space $S$ is a quasiprojective complex algebraic variety; the characterization is given in terms of the non-vanishing of a power of the invariant Green current associated to the family of maps. This conjectural characterization is inspired by and generalizes the Relative Manin-Mumford Conjecture for families of abelian varieties, recently proved by Gao and Habegger, and it includes as special cases the Manin-Mumford Conjecture (theorem of Raynaud) and the Dynamical Manin-Mumford Conjecture (posed by Ghioca, Tucker, and Zhang). We provide examples where the equivalence is known to hold, and we show that several recent results can be viewed as special cases. Finally, we give the proof of one implication in the conjectural characterization.
1. Introduction
Let $S$ be a smooth and irreducible quasiprojective variety defined over the field $\mathbb{C}$ of complex numbers. Fix an integer $d\geq 2$. An algebraic family of endomorphisms of $\mathbb{P}^{N}$ of degree $d$ is a morphism
$\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$
given by $\Phi(s,z)=(s,f_{s}(z))$, where each $f_{s}$ is an endomorphism of the complex projective space $\mathbb{P}^{N}$ of degree $d$. Throughout this note, we let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ denote a closed irreducible subvariety which is flat over $S$. We will use boldface $\mathbf{X}$ to denote the generic fiber of $\mathcal{X}$ and let $\mathbf{\Phi}:\mathbf{P^{N}}\to\mathbf{P^{N}}$ be the map induced by $\Phi$, viewed as an endomorphism over the function field $\mathbb{C}(S)$.
Inspired by Pink’s conjectures [Pi1] (especially [Pi2, Conjecture 6.2]), the recent theorems of Gao and Habegger on families of abelian varieties [GH2, Theorems 1.1 and 1.3], and by Zhang’s proposed extensions to more general dynamical systems on projective varieties [Zh], we discuss the following aim: to characterize the subvarieties $\mathcal{X}\subset S\times\mathbb{P}^{N}$ which contain a Zariski-dense set of preperiodic points of $\Phi$.
This is an ambitious goal; even the case where $S$ is a point remains vastly open, where a conjectural characterization goes by the name of the Dynamical Manin-Mumford Conjecture, proposed by Shouwu Zhang [Zh] and reformulated in [GTZ] and [GT]. Roughly speaking, if $\dim S=0$, the subvariety $\mathcal{X}$ of $\mathbb{P}^{N}$ should contain a Zariski-dense set of preperiodic points if and only if it is itself preperiodic (or is preperiodic for an endomorphism that commutes with $\Phi$). For $\dim S>0$, we will see that the variety $\mathcal{X}$ needs only be “big enough” in a preperiodic subvariety.
To make a precise conjecture, we employ the notion of $\Phi$-special subvarieties of Ghioca and Tucker in [GT]. We say that an irreducible subvariety $\mathcal{Y}\subset S\times\mathbb{P}^{N}$, which is flat over a Zariski-open subset of $S$ and Zariski-closed in $S\times\mathbb{P}^{N}$, is $\Phi$-special if there exist a subvariety $\mathbf{Z}\subset\mathbf{P^{N}}$ over the algebraic closure $\overline{\mathbb{C}(S)}$ containing the generic fiber ${\bf Y}$, a polarizable endomorphism $\mathbf{\Psi}:\mathbf{Z}\to\mathbf{Z}$, and an integer $n\in\mathbb{N}$ so that the following hold:
  • $\mathbf{\Phi}^{n}(\mathbf{Z})=\mathbf{Z}$;
  • $\mathbf{\Phi}^{n}\circ\mathbf{\Psi}=\mathbf{\Psi}\circ\mathbf{\Phi}^{n}$ on $\mathbf{Z}$; and
  • $\mathbf{Y}$ is preperiodic for $\mathbf{\Psi}$.
Recall that an endomorphism $\Psi$ of a projective variety $Z$ is polarizable if there exist an ample line bundle $L$ on $Z$ and integer $q>1$ so that $\Psi^{*}L\simeq L^{q}$. Polarizable is equivalent to being the restriction of an endomorphism $\tilde{\Psi}:\mathbb{P}^{N}\to\mathbb{P}^{N}$ to an invariant subvariety, where the embedding $Z\hookrightarrow\mathbb{P}^{N}$ may be defined by a power of $L$ and $\Psi=\tilde{\Psi}|_{Z}$; see, for example, [Fa1] [MZ1] [MZ2] for more on polarized endomorphisms.
Let $r_{\Phi,\mathcal{X}}$ denote the relative special dimension of $\mathcal{X}$ over $S$; that is,
$r_{\Phi,\mathcal{X}}:=\min\{\dim_{S}\mathcal{Y}\;:\;\mathcal{Y}~{}\text{ is }\Phi\text{-special and }\mathcal{X}\subset\mathcal{Y}\}$
is the minimal relative dimension over $S$ of a $\Phi$-special subvariety containing $\mathcal{X}$, where
$\dim_{S}\mathcal{Y}=\dim\mathcal{Y}-\dim S$
is the dimension of a general fiber of the projection to $S$. Note that $0\leq\dim_{S}\mathcal{X}\leq r_{\Phi,\mathcal{X}}\leq N$. We remark that there is not necessarily a “smallest $\Phi$-special subvariety containing $\mathcal{X}$”, in contrast with the setting of abelian varieties, as the intersection of $\Phi$-special subvarieties is not necessarily $\Phi$-special; see [GT, Example 3.1] for an example when $\dim S=0$.
Conjecture 1.1.
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree $>1$, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be a complex, irreducible subvariety which is flat over $S$. The following are equivalent.
  1. (1)
    $\mathcal{X}$ contains a Zariski-dense set of $\Phi$-preperiodic points.
  2. (2)
    $\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}\wedge[\mathcal{X}]\not=0$ for the relative special dimension $r_{\Phi,\mathcal{X}}$.
Here $\hat{T}_{\Phi}$ is the canonical Green current associated to $\Phi$ on $S\times\mathbb{P}^{N}$. We briefly recall its construction, as given for instance in [GV, §2.3]. Let $\omega$ be the Fubini-Study form on $\mathbb{P}^{N}$, and let $\hat{\omega}$ be the smooth $(1,1)$-form on $S\times\mathbb{P}^{N}$ defined by pulling back $\omega$ via the projection $S\times\mathbb{P}^{N}\to\mathbb{P}^{N}$. Then the sequence $d^{-n}(\Phi^{n})^{*}(\hat{\omega})$ converges weakly to the closed positive $(1,1)$-current $\hat{T}_{\Phi}$ on $S\times\mathbb{P}^{N}$; the potentials converge locally uniformly. If $\mathcal{Y}\subset S\times\mathbb{P}^{N}$ is a $\Phi$-special subvariety which is invariant for an endomorphism $\Psi$ that commutes with an iterate of $\Phi$, then $\hat{T}_{\Phi}=\hat{T}_{\Psi}$ on $\mathcal{Y}$ because of the commuting relation. Moreover, the slices of $\hat{T}_{\Phi}^{\wedge\dim_{S}\mathcal{Y}}$ in fibers of $\mathcal{Y}\to S$ are the measures of maximal entropy for the restriction of $\Psi$; see, for example, [DS2]. So condition (2) means that (an iterate of) $\mathcal{X}$ intersects these families of measures nontrivially. By convention, $\hat{T}^{\wedge 0}=1$. We point out here that the non-vanishing of $\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}\wedge[\mathcal{X}]$ can be thought of from an arithmetic viewpoint in view of the theory of adelic metrized line bundles on quasi-projective varieties developed by Yuan–Zhang [YZ1]. In the notation of [Guo] and [YZ1], it is equivalent to the non-vanishing of an intersection number $\tilde{L_{\Phi}}^{\wedge r_{\Phi,\mathcal{X}}}|_{X}\cdot\overline{H}^{\dim\mathcal{X}-r_{\Phi,\mathcal{X}}}$ for some $\overline{H}\in\widetilde{\mathrm{Pic}}(X/\mathbb{C})$ as shown in [Guo, Theorem 1.2].
Conjecture 1.1 is a generalization of the recent theorem of Gao and Habegger on families of abelian varieties [GH2, Theorem 1.3], when $\Phi$ preserves an abelian scheme $\mathcal{A}$ over $S$, inducing a homomorphism, and $\mathcal{X}$ is contained in $\mathcal{A}$. The preperiodic points of $\Phi$ in $\mathcal{A}$ coincide with the torsion points for the group structure. The current $\hat{T}_{\Phi}$ restricts to a Betti form on $\mathcal{A}$, and condition (2) of Conjecture 1.1 is equivalent to saying that $\mathcal{X}$ has maximal Betti rank; see Section 2 for details.
Conjecture 1.1 is known to hold in all cases when $N=1$, as we explain in Section 3. Conjecture 1.1 is also known to hold in some settings when $\dim S=0$ and $N>1$, moving beyond the cases where $\Phi$ preserves an abelian variety; see §3.1. Moreover, in the following sections, we explain that many existing works and conjectures can be viewed as special cases of Conjecture 1.1.
A weaker form of Conjecture 1.1 can be formulated as follows:
Conjecture 1.2.
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be a family of morphisms of degree $>1$, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible, flat family of subvarieties over $S$ containing a Zariski-dense set of preperiodic points of $\Phi$. Then $\mathcal{X}$ has codimension $\leq\dim S$ in a $\Phi$-special subvariety.
Conjecture 1.2 is easily obtained from Conjecture 1.1, because $\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}\wedge[\mathcal{X}]\not=0$ implies that $\dim\mathcal{X}\geq r_{{\Phi,\mathcal{X}}}=\dim\mathcal{Y}-\dim S$ for a $\Phi$-special subvariety $\mathcal{Y}$ containing $\mathcal{X}$. A special case of Conjecture 1.2 in a dynamical setting was proposed as [MS, Conjecture 1.9]; see also [MS, Theorem 1.8].
Gao and Habegger pointed out in [GH2] that the converse assertion to Conjecture 1.2 is not true, with explicit examples and constructions in the setting where $\Phi$ fixes an abelian subscheme of $S\times\mathbb{P}^{N}$; see also [Gao]. This led them to formulate their [GH2, Theorem 1.3], which is what our Conjecture 1.1 aims to extend. Some of those counterexample constructions arose already in the (conjectural) characterization of subvarieties with geometric canonical height 0, in the Geometric Dynamical Bogomolov Conjecture formulated by Gauthier and Vigny [GV]. The conjecture in [GV] aims to extend theorems for abelian varieties over function fields of characteristic 0, proved in [GH1, CGHX]. We discuss the relation between [GV, Conjecture 1.9] and Conjecture 1.1 in Section 5.
We complete this note with a proof of the following implication.
Theorem 1.3.
Condition (2) implies condition (1) of Conjecture 1.1.
The proof of Theorem 1.3 is based on the methods of Dujardin [Du], Berteloot-Bianchi-Dupont [BBD], and Gauthier [Ga1] to study supports of bifurcation currents and measures. For endomorphisms of abelian varieties, the implication (2) $\implies$ (1) was straightforward and observed in [ACZ], as we explain in Section 2.
Remark 1.4.
In the case that both $\Phi$ and $\mathcal{X}$ are defined over $\overline{\mathbb{Q}}$, we expect Conjectures 1.1 and 1.2 to remain true upon replacing $\Phi$-preperiodic points by $\Phi$-small points. Here a sequence of points $\{x_{n}\}_{n}\subset\mathcal{X}(\overline{\mathbb{Q}})$ is called $\Phi$-small if $\hat{h}_{\Phi}(x_{n})\to 0$, where $\hat{h}_{\Phi}(x):=\hat{h}_{\Phi_{\pi(x)}}(x)$ is the fiber-wise Call–Silverman canonical height introduced in [CS], and $\pi:S\times\mathbb{P}^{N}\to S$ is the projection.
Contents of this article. We begin by introducing a dynamical notion of $\Phi$-rank of $\mathcal{X}$, and we compare Conjecture 1.1 to the theorems of Gao and Habbeger from [GH2]. In particular, Conjecture 1.1 characterizes subvarieties $\mathcal{X}\subset S\times\mathbb{P}^{N}$ of maximal $\Phi$-rank; see Conjecture 2.3. We also compare the conditions of Conjecture 1.1 to the concept of non-degeneracy of subvarieties $\mathcal{X}$, which was originally introduced for families of abelian varieties. We then provide a brief survey of known cases of Conjecture 1.1, outside of the setting of abelian schemes. In particular, we remind the reader of the Dynamical Manin-Mumford Conjecture, first posed in [Zh], which corresponds to Conjecture 1.1 in the case where $\dim S=0$. We show that Conjecture 1.1 is known to hold in dimension $N=1$ over any base $S$ (and is in fact equivalent to [De3, Theorem 1.1]), and we relate it to the concept of $J$-stability for maps on $\mathbb{P}^{1}$. We then show that a conjecture of [BD2], and the recent classification of “special curves” in the moduli space $\mathrm{M}_{d}^{1}$ of maps on $\mathbb{P}^{1}$ in [JX], is a special case of Conjecture 1.1; see Theorem 3.5. Moreover, we illustrate the strength of Conjecture 1.1 with an example showing that uniform versions of the conjecture are consequences of the conjecture itself; the example we provide comes from the study of shared preperiodic points for distinct maps on $\mathbb{P}^{1}$. For maps on $\mathbb{P}^{N}$, with $N>1$, we show that Conjecture 1.1 implies the recent sparsity theorem of Gauthier, Taflin, and Vigny in [GTV] about PCF maps in the moduli space $\mathrm{M}_{d}^{N}$ of maps on $\mathbb{P}^{N}$. We explore the concept of minimal $\Phi$-rank and compare Conjecture 1.1 to the Geometric Bogomolov Conjecture posed in [GV]. Finally, in Section 6, we provide the proof of Theorem 1.3.
Acknowledgements. We would like to thank Ziyang Gao, Thomas Gauthier, Lars Kühne, Harry Schmidt, and Gabriel Vigny for many interesting discussions and their help during the preparation of this note. We also thank the Simons Foundation for their support during a Symposium in August 2022 where this work was initiated. This project was supported in part with funding from the National Science Foundation and the Radcliffe Institute for Advanced Study.
2. $\Phi$-rank and Betti-rank
In this section, we compare Conjecture 1.1 with the theorem of Gao and Habbeger [GH2, Theorem 1.3], stated below as Theorem 2.4. In their setting of abelian varieties, the case of $\dim S=0$ reduces to the original Manin-Mumford Conjecture proved by Raynaud [Ra]. We also introduce a dynamical notion of rank that extends the notion of Betti rank from [ACZ]. Conjecture 2.3, which is equivalent to Conjecture 1.1, is a characterization of subvarieties of $S\times\mathbb{P}^{N}$ with maximal $\Phi$-rank.
2.1. $\Phi$-Rank
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree $>1$, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible flat subvariety over $S$. We define the $\Phi$-rank of $\mathcal{X}$ to be
$\operatorname{rank}_{\Phi}(\mathcal{X}):=\max\{r\geq 0~{}:~{}\hat{T}_{\Phi}^{\wedge r}\wedge[\mathcal{X}]\not=0\}.$
It is clear from the definition that $\operatorname{rank}_{\Phi}(\mathcal{X})\leq\dim\mathcal{X}$.
Lemma 2.1.
For $\dim S=0$, so that $\Phi$ is a single endomorphism $f:\mathbb{P}^{N}\to\mathbb{P}^{N}$, and for any irreducible subvariety $Z$ of $\mathbb{P}^{N}$, we have
$\operatorname{rank}_{f}(Z)=\dim Z.$
Proof.
Let $\ell=\dim Z$. If $\ell=0$, then the statement is clear. So we may assume that $\ell>0$. Let $\omega$ be the Fubini-Study form on $\mathbb{P}^{N}$, representing $c_{1}(\mathcal{O}(1))$. If $f$ has degree $d$, then $d^{-n}(f^{n})^{*}\omega$ is cohomologous to $\omega$ for all $n\geq 1$. In particular, we have
[…]
for all $n\geq 1$. Because of the local-uniform convergence of the potentials of pullbacks of $\omega$ to that of the invariant current $T_{f}$, we have
$\left(d^{-\ell n}(f^{n})^{*}\omega^{\wedge\ell}\right)\wedge[Z]\to T_{f}^{\wedge\ell}\wedge[Z]$
as $n\to\infty$ [Dem, Chapter III Corollary 3.6], allowing us to conclude that $T_{f}^{\wedge\ell}\wedge[Z]\not=0$. On the other hand, we have $T_{f}^{\wedge(\ell+1)}\wedge[Z]=0$ for dimension reasons. ∎
Proposition 2.2.
If $\mathcal{Y}\subset S\times\mathbb{P}^{N}$ is a $\Phi$-special subvariety, then
$\hat{T}_{\Phi}^{\wedge(1+\dim_{S}\mathcal{Y})}\wedge[\mathcal{Y}]=0.$
In particular, for any irreducible $\mathcal{X}\subset S\times\mathbb{P}^{N}$ which is flat over $S$, we have
$\dim_{S}\mathcal{X}\leq\operatorname{rank}_{\Phi}(\mathcal{X})\leq r_{\Phi,\mathcal{X}}$
for the relative special dimension $r_{\Phi,\mathcal{X}}$ of $\mathcal{X}$.
Proof.
Suppose that $\mathcal{Y}$ is $\Phi$-special, and set $k=\dim_{S}\mathcal{Y}$. Let $\Psi$ be a family of polarized endomorphisms over $S$ that commutes with an iterate of $\Phi$ along $\mathcal{Z}$ with $\mathcal{Y}\subset\mathcal{Z}\subset S\times\mathbb{P}^{N}$; replacing $\Psi$ with an iterate, we assume that $\mathcal{Y}$ is prefixed for $\Psi$. Now replace $\mathcal{Y}$ with an iterate under $\Psi$ so that $\Psi(\mathcal{Y})=\mathcal{Y}$, and pass to a normalization if it is not itself normal. We have $\hat{T}_{\Phi}=\hat{T}_{\Psi}$ on $\mathcal{Y}$. Since $(\Psi,\mathcal{Y})$ defines a family of polarized dynamical systems of degree $e>1$, there is a (1,1)-form $\omega$ on $\mathcal{Y}$ representing the first Chern class of the polarizing (relatively ample) line bundle $\mathcal{L}$ so that $\omega^{k+1}=0$. Because of the local-uniform convergence of the potentials of pullbacks of $\omega$ to a potential for $\hat{T}_{\Psi}$, we know that the wedge power $\hat{T}_{\Psi}^{\wedge j}$ is the limit of pullbacks $\frac{1}{e^{jn}}(\Psi^{n})^{*}(\omega^{\wedge j})$ on $\mathcal{Y}$, for any $j\geq 1$. As $\omega^{\wedge(k+1)}=0$, we have $\hat{T}_{\Phi}^{\wedge(1+\dim_{S}\mathcal{Y})}\wedge[\mathcal{Y}]=0$.
The upper bound on $\operatorname{rank}_{\Phi}(\mathcal{X})$ is an immediate consequence, because $\mathcal{X}$ is contained in a $\Phi$-special subvariety of relative dimension $r_{\Phi,\mathcal{X}}$.
For the lower bound on $\operatorname{rank}_{\Phi}(\mathcal{X})$, observe that the slice $(1,1)$-current $T_{s}$ of $\hat{T}_{\Phi}$ over $s\in S$ satisfies $T_{s}^{\wedge\ell}\wedge[X]\not=0$ for every subvariety $X$ of dimension $\ell$ in $\mathbb{P}^{N}$ and every $s\in S$, by Lemma 2.1. The nonvanishing of $\hat{T}_{\Phi}^{\wedge\dim_{S}\mathcal{X}}\wedge[\mathcal{X}]$ follows; see, for example, [BB, Proposition 4.3] on slicing. ∎
By Proposition 2.2, Conjecture 1.1 becomes a characterization of subvarieties $\mathcal{X}$ of maximal $\Phi$-rank, in the following sense:
Conjecture 2.3.
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree $>1$, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible subvariety which is flat over $S$. The following are equivalent:
  1. (1)
    $\mathcal{X}$ contains a Zariski dense set of $\Phi$-preperiodic points.
  2. (2)
    $\operatorname{rank}_{\Phi}(\mathcal{X})=r_{\Phi,\mathcal{X}}$.
2.2. Families of abelian varieties
Let us now compare Conjecture 2.3 with the following theorem of Gao and Habegger.
Theorem 2.4.
[GH2, Theorem 1.3] Let $\mathcal{A}\to S$ be an abelian scheme of relative dimension $g\geq 1$. Assume that $\mathcal{X}\subset\mathcal{A}$ is a closed irreducible subvariety, flat over $S$, for which the orbit $\mathbb{Z}\cdot\mathcal{X}$ is Zariski dense in $\mathcal{A}$. The following are equivalent:
  1. (1)
    $\mathcal{X}$ contains a Zariski dense set of torsion points in $\mathcal{A}$.
  2. (2)
    $\operatorname{rank}_{\mathrm{Betti}}(\mathcal{X})=2g$.
Here $\operatorname{rank}_{\mathrm{Betti}}(\mathcal{X})$ is the generic Betti rank of $\mathcal{X}$. Its study was initiated by André-Corvaja-Zannier [ACZ] and it has now become ubiquitous in issues related to unlikely intersections. We briefly recall its definition and then explain why Conjecture 2.3 is a generalization of Theorem 2.4.
Let $\pi:\mathcal{A}\to S$ denote the projection, and let $\Delta\subset S$ be a simply connected open subset. Choose holomorphic functions $\omega_{i}:\Delta\to\mathbb{C}^{g}$, $i=1,\ldots,2g$, defining a basis of the period lattice of the fibers, so that
$\mathcal{A}_{s}\simeq\mathbb{C}^{g}/\omega_{1}(s)\mathbb{Z}\oplus\cdots\oplus\omega_{2g}(s)\mathbb{Z}$
for $s\in\Delta$, with the isomorphism denoted by
$x\mapsto\sum_{i=1}^{2g}\beta_{i}(x)\omega_{i}(s)$
for $\beta_{i}(x)\in\mathbb{R}/\mathbb{Z}$. The Betti map $b_{\Delta}:\pi^{-1}(\Delta)\to\mathbb{R}^{2g}/\mathbb{Z}^{2g}$ is a real-analytic map defined by
$b_{\Delta}(x)=(\beta_{1}(x),\ldots,\beta_{2g}(x))$
for $x\in\mathcal{A}_{s}$. The generic Betti rank $\operatorname{rank}_{\mathrm{Betti}}(\mathcal{X})$ is the maximal rank of the differential of the Betti map over all points in $\mathcal{X}\cap\pi^{-1}(\Delta)$, and it is independent of the choice of simply connected $\Delta$. Associated to the Betti map and a choice of polarization $\mathcal{L}$ is a $(1,1)$-form $\omega_{\mathcal{A},\mathcal{L}}$ on $\mathcal{A}$, called the Betti form and first introduced by Mok [Mo]; see [DGH2, Proposition 2.2]. It can be defined by
$\omega_{\mathcal{A},\mathcal{L}}=\sum_{1\leq i<j\leq 2g}a_{ij}\,d\beta_{i}\wedge d\beta_{j},$
for constants $a_{ij}\in\mathbb{R}$; its restriction to each fiber $\mathcal{A}_{s}$ is the Riemann form associated to the choice of polarization. See, for example, [CGHX, §2.4].
To see that Conjecture 2.3 is a generalization of Theorem 2.4, let us first replace $S$ with a Zariski-open subset so that all fibers of $\mathcal{A}\to S$ are (smooth) abelian varieties. Consider the multiplication-by-$M$ morphism
$[M]:\mathcal{A}\to\mathcal{A}$
for any choice of integer $M\geq 2$. This endomorphism is polarizable with degree $M^{2}$, and thus induces an embedding
$\mathcal{A}\hookrightarrow S\times\mathbb{P}^{N}$
for some $N$, so that, as in [Fa1, Corollary 2.2], $[M]$ extends to an algebraic family of morphisms
$\Phi_{M,\mathcal{A}}:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}.$
The Green current $\hat{T}_{\Phi_{M,\mathcal{A}}}$ on $S\times\mathbb{P}^{N}$ restricts to the Betti form $\omega_{\mathcal{A},\mathcal{L}}$ on $\mathcal{A}$ associated to the polarization $\mathcal{L}$; indeed, the current is uniquely determined (in its cohomology class $c_{1}(\mathcal{L})$) by its invariance under pullback for the restriction $\Phi_{M,\mathcal{A}}|\mathcal{A}$, where $[M]^{*}\omega_{\mathcal{A},\mathcal{L}}=M^{2}\omega_{\mathcal{A},\mathcal{L}}$. It is then an immediate consequence of [DGH2, Proposition 2.2 (iii)] that
$\operatorname{rank}_{\mathrm{Betti}}(\mathcal{X})=2\operatorname{rank}_{\Phi_{M,\mathcal{A}}}(\mathcal{X}).$
It remains to relate the hypothesis on $\mathcal{X}$, that its $\mathbb{Z}$-orbit is Zariski dense in $\mathcal{A}$, to our notion of relative special dimension. Indeed, the condition on $\mathcal{X}\subset\mathcal{A}$ implies that $\mathcal{X}$ is not contained in any proper subgroup scheme over $S$ nor a torsion-translate of such a subgroup. In particular, the smallest $\Phi_{M,\mathcal{A}}$-special subvariety containing $\mathcal{X}$ in $S\times\mathbb{P}^{N}$ is the embedded copy of $\mathcal{A}$ itself. That is, $r_{\Phi,\mathcal{X}}=g$, the relative dimension of $\mathcal{A}$.
Remark 2.5.
Comparing the statement of Theorem 2.4 with Conjectures 1.1 and 2.3, it is important to notice some additional complications that arise in the dynamcial setting. It can happen that the $\Phi$-orbit of a subvariety $\mathcal{X}\subset S\times\mathbb{P}^{N}$ is Zariski dense in some $\Phi$-special subvariety $\mathcal{Y}\subset S\times\mathbb{P}^{N}$, while $r_{\Phi,\mathcal{X}}<\dim_{S}\mathcal{Y}$. Such examples arose in the formulation of the Dynamical Manin-Mumford conjecture; see [GTZ] for details. Those examples came from abelian varieties, where the $\Phi$ in question is built from complex multiplication, while the smaller $\Phi$-special is invariant for the usual $\mathbb{Z}$-action.
Remark 2.6.
The statement of Theorem 1.3, that (2) implies (1) in Conjecture 1.1, is easy to prove in the setting of abelian varieties, due to the smoothness and analytic properties of the Betti forms [ACZ, Proposition 2.1.1]. (The proof in [ACZ] is written for sections of a family $\mathcal{A}\to S$ of abelian varieties; to treat the case of a subvariety $\mathcal{X}$ in $\mathcal{A}$, notice that $\mathcal{X}$ is the image of the identity section of the base extension $\mathcal{A}\times_{S}\mathcal{X}\to\mathcal{X}$.)
2.3. Non-degeneracy
We conclude this section by observing that there is an important notion in the literature which is similar to but distinct from condition (2) of Conjecture 1.1; we include it here for comparison. By definition, $\mathcal{X}$ in $S\times\mathbb{P}^{N}$ is non-degenerate for $\Phi$ if $\hat{T}_{\Phi}^{\wedge\dim\mathcal{X}}\wedge[\mathcal{X}]\not=0$. This concept was introduced for subvarieties $\mathcal{X}$ in families of powers of elliptic curves by Habegger [Ha] and played an important role in the proof of the Geometric Bogomolov Conjecture over characteristic 0 by Gao–Habegger and Cantat–Gao–Habegger–Xie [GH1, CGHX], the proof of the uniform Mordell–Lang conjecture by Gao–Habegger, Kühne and Gao–Ge–Kühne [DGH1, DGH2, , GGK] and the proof of the relative Manin-Mumford conjecture by Gao–Habegger [GH2]. In [YZ1], this notion of non-degeneracy was extended to families of polarized dynamical systems, and it is a key hypothesis for their theorems on arithmetic equidistribution. In particular, if $\Phi$ and $\mathcal{X}$ are defined over $\overline{\mathbb{Q}}$, then Conjecture 1.1 would characterize the existence of a Zariski-dense set of small points in $\mathcal{X}$, while the non-degeneracy condition would imply that the Galois orbits of these small points are uniformly distributed with respect to the measure $\hat{T}_{\Phi}^{\wedge\dim\mathcal{X}}\wedge[\mathcal{X}]\not=0$ (if the points exist) [YZ1, Theorem 5.4.3] [Ga2]. Note that every projective algebraic subvariety of $\mathbb{P}^{N}$ is non-degenerate for a morphism $f:\mathbb{P}^{N}\to\mathbb{P}^{N}$ when $\dim S=0$, by Lemma 2.3.
3. Some known cases of Conjecture 1.1
In this section, we provide examples where Conjecture 1.1 is known, moving away from the setting of abelian varieties. We begin with the original Dynamical Manin-Mumford Conjecture, corresponding to the case of $\dim S=0$. In §3.2, we observe that the Conjecture holds in all cases when $N=1$, as it can be deduced from [De3, Theorem 1.1], and we relate it to the theory of $J$-stability for maps on $\mathbb{P}^{1}$. In §3.3 we explain that Conjecture 1.1 implies the so-called Dynamical André-Oort conjecture (or “DAO”). Finally, we illustrate that uniform versions of the conjecture are consequences of the conjecture itself, with the example of shared preperiodic points for distinct maps on $\mathbb{P}^{1}$.
3.1. The Dynamical Manin-Mumford Conjecture
When $S$ is a point, so that $\dim S=0$ and $\Phi:\mathbb{P}^{N}\to\mathbb{P}^{N}$ is an endomorphism over $\mathbb{C}$, Conjectures 1.1 and 1.2 reduce to Zhang’s Dynamical Manin-Mumford Conjecture (or DMM), as reformulated by Ghioca and Tucker in [GT]: a subvariety $X$ of $\mathbb{P}^{N}$ contains a Zariski-dense set of preperiodic points if and only if it is a $\Phi$-special subvariety. Indeed, the current $T_{\Phi}^{\wedge r}\wedge[X]$ will vanish if $r>\dim X$, but is nonzero for $r=\dim X$ by Lemma 2.1. The implication that a $\Phi$-special subvariety of $\mathbb{P}^{N}$ always contains a dense set of preperiodic points is well known; see for example [Fa1, BD, DS1]. The converse implication has been proved in few settings outside of the cases of endomorphisms of abelian varieties.
It is worth observing that DMM is obvious for maps on $\mathbb{P}^{1}$: a single point is either preperiodic (which is equivalent to special for points) or it is not. The DMM conjecture has been fully resolved for polarized endomorphisms of $(\mathbb{P}^{1})^{N}$ over $\mathbb{C}$ [GNY, GNY, MSW], for subvarieties $X\subset(\mathbb{P}^{1})^{N}$ of arbitrary dimension. Very recently, Dujardin, Favre, and Ruggiero completed a proof for polynomial maps of $\mathbb{A}^{2}$ that extend regularly to $\mathbb{P}^{2}$, where $X$ is a complex algebraic curve in $\mathbb{P}^{2}$ (though there remains a condition on the curve, regarding its intersection with the line at infinity) [DFR].
A related problem for polynomial automorphisms of $\mathbb{A}^{2}$ was treated in [DF]. We do not consider their setting in this article.
3.2. Dimension $N=1$
Here we show that Conjecture 1.1 in the case of $N=1$ follows from [De3, Theorem 1.1].
Theorem 3.1.
Conjecture 1.1 holds in dimension $N=1$.
Proof.
An irreducible flat subvariety $\mathcal{X}$ of $S\times\mathbb{P}^{1}$, if not equal to all of $S\times\mathbb{P}^{1}$, is a multi-section over $S$. That is, after replacing the parameter space $S$ with a branched cover, we may assume that $\mathcal{X}$ is the graph of a marked point $a:S\to\mathbb{P}^{1}$. If $r_{\Phi,\mathcal{X}}=0$, then $\mathcal{X}$ is itself $\Phi$-special, meaning that the marked point $a$ is persistently preperiodic for $\Phi$. In this case, the preperiodic points are obviously dense in $\mathcal{X}$ and the current $\hat{T}_{\Phi}^{0}\wedge[\mathcal{X}]=[\mathcal{X}]$ is clearly nonzero, so the equivalence of (1) and (2) in Conjecture 1.1 holds. If $r_{\Phi,\mathcal{X}}=1$, then the current of (2) is nonzero if and only if the point is unstable, in the sense of [De3]. In other words, the sequence of holomorphic maps $\{s\mapsto\Phi_{s}^{n}(a(s))\}$ fails to formal a normal family on the parameter space $S$; see, for example, [De2, Theorem 9.1] for the equivalence of normality and the vanishing of $\hat{T}_{\Phi}\wedge[\mathcal{X}]$. As proved in [De3, Theorem 1.1], if the point $a$ is not persistently preperiodic, then stability on all of $S$ implies that the family $\Phi$ is isotrivial and the point $a$ will never be preperiodic. On the other hand, instability implies, by Montel’s theory of normal families, that the point $a$ will be preperiodic for a Zariski-dense set of parameters; see, for example, [De3, Proposition 5.1]. So the equivalence in Conjecture 1.1 holds also for $r_{\Phi,\mathcal{X}}=1$. ∎
Remark 3.2.
Conjecture 1.1 also implies [De3, Theorem 1.1] (and therefore also [DF, Theorem 2.5] and [Mc, Lemma 2.1] addressing the case of marked critical points in $\mathbb{P}^{1}$), so it is logically equivalent in dimension $N=1$. Indeed, suppose that $\Phi:S\times\mathbb{P}^{1}\to S\times\mathbb{P}^{1}$ is an algebraic family of maps of degree $d>1$, and suppose that $\Gamma_{a}\subset S\times\mathbb{P}^{1}$ is the graph of a marked point $a:S\to\mathbb{P}^{1}$. Assume the pair $(\Phi,a)$ is stable, meaning that $\hat{T}_{\Phi}\wedge[\Gamma_{a}]=0$. We will deduce from Conjecture 1.1 that $a$ is either persistently preperiodic or the pair $(\Phi,a)$ is isotrivial. It suffices to assume that $\dim S=1$. Passing to a branched cover of $S$ if necessary, we can mark three distinct periodic points for $\Phi$ and, removing a Zariski-closed subset of $S$ where they collide, we can change coordinates on $\mathbb{P}^{1}$ so that $\{0,1,\infty\}$ are persistently periodic. Stability of $(\Phi,a)$ implies one of two things: either (1) the graph $\Gamma_{a}$ is itself preperiodic (i.e., the special dimension is $r_{\Phi,\Gamma_{a}}=0$), or (2) as a consequence of Conjecture 1.1 with $r_{\Phi,\Gamma_{a}}=1$, the preperiodic points of $\Phi$ cannot be Zariski dense in $\Gamma_{a}$. In case (1), we are done. In case (2), this means that there is a Zariski-open subset $U$ of $S$ over which the point $a$ is never preperiodic. In particular, the point $a$ and its infinite forward orbit is disjoint from the set $\{0,1,\infty\}$ over the quasiprojective curve $U$. But, as McMullen observed in his proof of [Mc, Lemma 2.1], there are at most finitely many non-constant holomorphic functions $U\to\mathbb{P}^{1}\setminus\{0,1,\infty\}$. This implies that the iterates $\Phi^{n}(a)$ must be constant (and distinct) over $U$, for all $n$ sufficiently large. But then, by a simple interpolation argument, we see that $\Phi_{s}$ is independent of the parameter $s\in U$. In other words, in case (2), the family $\Phi$ is isotrivial.
Remark 3.3.
Suppose that $\Phi:S\times\mathbb{P}^{1}\to S\times\mathbb{P}^{1}$ is an algebraic family of endomorphisms on $\mathbb{P}^{1}$ and that
$\mathcal{X}=\mathrm{Crit}(\Phi)$
is the critical locus. (This $\mathcal{X}$ is not necessarily irreducible, but we can apply Theorem 3.1 to each component.) Then $r_{\Phi,\mathcal{X}}=0$ (for all components) if and only if $\Phi$ is a family of postcritically finite maps. We know from Thurston rigidity that this can hold if and only if $\Phi$ is either an isotrivial family or a family of flexible Lattès maps [DH].
For $r_{\Phi,\mathcal{X}}=1$, the current of Conjecture 1.1 condition (2),
$\hat{T}_{\Phi}\wedge[\mathrm{Crit}(\Phi)],$
projects to the parameter space $S$ as the bifurcation current $T_{\mathrm{bif}}$ for the family $\Phi$. See [De1] and [DF] for definitions. It was proved in [De1] that the bifurcation current vanishes if and only if the family is $J$-stable in the sense of Mañé-Sad-Sullivan [MSS] and Lyubich [Ly]. So Conjecture 1.1 (i.e., Theorem 3.1) implies the (known) characterization of instability in algebraic families $\Phi$ by the existence of many non-persistent preperiodic critical points; compare [Mc, Theorem 2.2] and [DF, Theorem 2.5].
We return to the topic of $J$-stability in higher dimensions in Section 5.
3.3. DAO as a special case
Let $f:S\times\mathbb{P}^{1}\to S\times\mathbb{P}^{1}$ be an algebraic family of rational maps on $\mathbb{P}^{1}$ of degree $d\geq 2$. We say it has dimension $m$ in moduli if the induced projection from $S$ to $\mathrm{M}_{d}^{1}$, the moduli space of all maps of degree $d$ on $\mathbb{P}^{1}$, has $m$-dimensional image. Here we observe that Conjecture 1.1 (in fact, in its weaker form of Conjecture 1.2) implies the following conjecture proposed by Baker-DeMarco [BD2] and Ghioca-Hsia-Tucker [GHT1]; the statement here is given explicitly in [De3, Conjecture 6.1].
Conjecture 3.4.
Assume that $\dim S=m>0$ and that $f$ has dimension $m$ in moduli. Let $a_{0},\ldots,a_{m}:S\to\mathbb{P}^{1}$ be $m+1$ marked points, and assume that there is a Zariski-dense set of parameters $s\in S$ such that $a_{0}(s),\ldots,a_{m}(s)$ are simultaneously $f_{s}$-preperiodic. Then the marked points $a_{0},\ldots,a_{m}$ are dynamically related along $S$.
By definition, marked points $(a_{0},\ldots,a_{m})$ are dynamically related along $S$ if there exists a proper subvariety $\mathcal{Y}\subset S\times(\mathbb{P}^{1})^{m+1}$ projecting dominantly to $S$ which is preperiodic for the fiber-product $\Phi=f^{[m+1]}$ defined by
(3.1) $\Phi(s,z_{0},\ldots,z_{m})=(s,f_{s}(z_{0}),\ldots,f_{s}(z_{m}))$
and which contains the graph $\Gamma$ of $(a_{0},\ldots,a_{m})$ over $S$.
The converse implication to Conjecture 3.4, that a dynamical relation implies the density of the simultaneous preperiodic points, was proved in [De3]. Special cases of Conjecture 3.4 and closely related results were obtained in [MZ1, MZ2, BD1, BD2, DWY1, DWY2, GHN, GHT1, GHT2, GHT3, GKN, GKNY, GY, FG1, FG2, JX].
We wish to emphasize one case of Conjecture 3.4: if a subvariety $V$ of the moduli space $\mathrm{M}_{d}^{1}$ contains a Zariski-dense set of postcritically finite maps, then the subvariety $V$ is conjectured to be “special”, meaning that every $(1+\dim V)$-tuple of critical points is dynamically related along $V$. One works with an algebraic family $\Phi$ over a parameter space $S$ that maps dominantly to $V$, with marked critical points $\{c_{1},\ldots,c_{2d-2}\}$. If $m=\dim V$, one considers all subsets of these critical points of size $m+1$. In this form, the conjecture was dubbed the “Dynamical André-Oort Conjecture” in [Si, Chapter 6], or “DAO” in [JX], because of parallels between the theory of postcritically finite maps and of elliptic curves with complex multiplication. This case of the conjecture was resolved for algebraic curves in $\mathrm{M}_{d}^{1}$ in [JX] but remains open for higher-dimensional subvarieties of $\mathrm{M}_{d}^{1}$.
Theorem 3.5.
Conjecture 1.2 implies Conjecture 3.4.
Proof.
To place Conjecture 3.4 in the setting of Conjectures 1.1 and 1.2, note that the $\Phi$ of (3.1) defines a family of polarized endomorphisms of $(\mathbb{P}^{1})^{m+1}$ and so extends to a family of endomorphisms of projective space $\mathbb{P}^{M}$ for some dimension $M\geq m+1$. In particular, the collection $(a_{0},\ldots,a_{m})$ is dynamically related if and only if the graph $\Gamma$ of this $(m+1)$-tuple lies in a $\Phi$-special subvariety of relative dimension $\leq m$ over $S$. Indeed, for powers of nonisotrivial maps on $\mathbb{P}^{1}$ such as $\Phi$, all $\Phi$-special subvarieties in $S\times(\mathbb{P}^{1})^{m+1}$ will be preperiodic for $\Phi$; see, for example, the discussion in Section 2 of [GT], as Lattès maps coming from complex multiplication are rigid in $\mathrm{M}_{d}^{1}$.
Note also that $\Gamma$ contains a Zariski dense set of $\Phi$-preperiodic points if and only if there is a Zariski dense set of $s\in S$ at which $a_{0}(s),\ldots,a_{N}(s)$ are simultaneously $f_{s}$-preperiodic.
So assume that Conjecture 1.2 holds for $\mathcal{X}=\Gamma$ and $\Phi=f^{[m+1]}$, and suppose that the hypotheses of Conjecture 3.4 hold. Then $\Gamma$ must have codimension $\leq m$ in a $\Phi$-special subvariety. As $m=\dim S=\dim\Gamma$, we must have $r_{\Phi,\Gamma}\leq m$, so that the graph $\Gamma$ must lie in a $\Phi$-special subvariety of relative dimension $\leq m$. In other words, the marked points are dynamically related. ∎
3.4. Common preperiodic points in $\mathbb{P}^{1}$ and uniform bounds
It was proved in [BD1] and [YZ2] that any pair of maps $f,g:\mathbb{P}^{1}\to\mathbb{P}^{1}$ over $\mathbb{C}$, of degrees $>1$, will satisfy either $\mathrm{Prep}(f)=\mathrm{Prep}(g)$ or $|\mathrm{Prep}(f)\cap\mathrm{Prep}(g)|<\infty$. This result can be viewed as a special case of Conjecture 1.1, at least when $\deg f=\deg g$. Indeed, suppose that $\deg(f)=\deg(g)=d\geq 2$, and consider the action of $\Phi:=(f,g)$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$. This $\Phi$ is polarizable and so extends to an endomorphism of some $\mathbb{P}^{N}$, restricting to the given map $(f,g)$ on an embedded copy of $\mathbb{P}^{1}\times\mathbb{P}^{1}$. The common preperiodic points in $\mathbb{P}^{1}$ for $f$ and $g$ correspond to preperiodic points of $\Phi$ in the diagonal $\Delta\subset\mathbb{P}^{1}\times\mathbb{P}^{1}$. Conjecture 1.1 reduces to the DMM conjecture in this setting, as discussed in §3.1; explicitly, it states that $\Delta$ is $\Phi$-special if and only if $|\mathrm{Prep}(f)\cap\mathrm{Prep}(g)|=\infty$. The equivalence of equality $\mathrm{Prep}(f)=\mathrm{Prep}(g)$ with $\Delta$ being $\Phi$-special follows from [YZ2, Theorem 1.4] combined with [LP, Theorem A] and [MS, Theorem 1.10].
Conjecture 1.1 also predicts the existence of uniform bounds on the size of $|\mathrm{Prep}(f)\cap\mathrm{Prep}(g)|$ in families of pairs. For example, it was proved in [DM] that there exists a Zariski-open and -dense subset $U_{d}$ of the space of all pairs $\mathrm{Rat}_{d}\times\mathrm{Rat}_{d}$ of degree $d$, and a constant $B_{d}$, so that
(3.2) $|\mathrm{Prep}(f)\cap\mathrm{Prep}(g)|\leq B_{d}$
for all $(f,g)\in U_{d}$; this theorem can be seen as a special case of Conjecture 1.1 (and in fact of Conjecture 1.2), as follows.
Fix a degree $d>1$ and let $S_{d}=\mathrm{Rat}_{d}\times\mathrm{Rat}_{d}$ be the space of all pairs of maps on $\mathbb{P}^{1}$ of degree $d$, and let
$\Phi:S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})\to S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})$
denote the corresponding algebraic family. For any integer $k\geq 1$, let $\Phi^{[k]}$ denote the $k$-th fiber power
$\Phi^{[k]}=(f,g,\ldots,f,g)$
on $S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$, so that $\Phi^{[1]}=\Phi$, and set
$\mathcal{X}_{k}=S_{d}\times\Delta^{k}$
in $S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$ for the diagonal $\Delta\subset\mathbb{P}^{1}\times\mathbb{P}^{1}$. Note that the codimension of $\mathcal{X}_{k}$ in $S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$ is equal to $k$, for all $k\geq 1$.
By [DM, Lemma 5.4], to prove (3.2) it suffices to show that there exists $k\geq 1$ so that $\mathcal{X}_{k}$ does not contain a Zariski dense set of $\Phi^{[k]}$-preperiodic points. We claim that the relative special dimension $r_{\Phi^{[k]},\mathcal{X}_{k}}$ is equal to $2k$ for any $k\geq 1$; that is, the subvariety $\mathcal{X}_{k}$ of $S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$ is not contained in any $\Phi^{[k]}$-special subvariety other than $S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$ itself. Then (3.2) follows immediately from Conjecture 1.2, for any $k>\dim S_{d}$.
Let $\mathcal{Z}\subset S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$ be ${\Phi}^{[k]}$-special and assume the generic fiber $\mathbf{Z}$ contains $\mathbf{X}_{k}$. Since there exists a pair of non-exceptional maps $(f_{s},g_{s})$ for $s\in S_{d}$ with distinct Julia sets, it must be that $\mathbf{Z}$ is $\mathbf{\Phi}^{[k]}$-preperiodic and neither $\mathbf{f}$ nor $\mathbf{g}$ is exceptional; this follows from the structural result on invariant subvarieties for product maps on $(\mathbb{P}^{1})^{N}$ (for any $N$) over fields of characteristic 0, proved in [MeSc] (and reproved by different methods in [GNY]). It follows from this classification that $\mathcal{Z}$ must be all of $S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$. This completes the proof.
4. Further consequences of Conjecture 1.1: sparsity of special subvarieties
For integers $N\geq 1$ and $d\geq 2$, we let $\mathrm{M}_{d}^{N}$ denote the affine complex-algebraic variety which is the space of conjugacy classes of endomorphisms on $\mathbb{P}^{N}$ with degree $d$ [Si]. In this section we show that the following conjecture is a consequence of Conjecture 1.1 (and in fact, of its weaker form, Conjecture 1.2). The case where $\mathcal{X}$ is the critical locus of $\Phi$ is a recent theorem of [GTV].
Conjecture 4.1.
Let $N>1$ and $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of endomorphisms of degree $d>1$ for which the induced map $S\to\mathrm{M}_{d}^{N}$ is dominant. Let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible flat family of subvarieties over $S$ that is neither all of $S\times\mathbb{P}^{N}$ nor has codimension $N$. Then the set of $s\in S(\mathbb{C})$ such that $X_{s}$ is $\Phi_{s}$-special is not Zariski dense in $S$.
Remark 4.2.
Conjecture 4.1 fails if we do not assume that $S\to\mathrm{M}_{d}^{N}$ is dominant. For example, $\Phi$ could be a family of regular polynomial maps on $\mathbb{P}^{2}$ (with coordinates $(x:y:z)$), where the line at infinity $\{z=0\}$ is invariant for $\Phi_{s}$, for all $s\in S$. Even if we further assume that $\mathcal{X}$ is not $\Phi$-special, the conjecture can fail without assuming dominance of $S\to\mathrm{M}_{d}^{N}$; for example, the family $\mathcal{X}$ of lines $\{x=s\,y\}$ in $\mathbb{P}^{2}$, for $s\in\mathbb{C}$, is not special under $\Phi(s,x,y)=(s,x^{2},y^{2})$ but becomes $\Phi_{s}$-preperiodic for each root of unity $s$.
Theorem 4.3.
Conjecture 1.2 implies Conjecture 4.1.
Proof.
If for some $s\in S(\mathbb{C})$ the subvariety $X_{s}\subset\mathbb{P}^{N}$ is $\Phi_{s}$-special, then $X_{s}$ must contain a Zariski dense set of preperiodic points (see the discussion in §3.1). It follows that, for any integer $m\geq 1$, if we consider the product map $\Phi_{s}^{[m]}=(\Phi_{s},\ldots,\Phi_{s})$ on $(\mathbb{P}^{N})^{m}$, then the $\Phi_{s}^{[m]}$-preperiodic points are also Zariski dense in the product $X_{s}^{m}$.
As for the example of §3.4, we consider fiber powers of the family $\Phi$. Let
$\Phi^{[m]}:S\times(\mathbb{P}^{N})^{m}\to S\times(\mathbb{P}^{N})^{m}$
denote the algebraic family formed by the products, for integers $m\geq 1$. Note that $\Phi^{[m]}$ is a polarizable endomorphism, so it extends as an algebraic family of morphisms
$\Phi_{N,m}:S\times\mathbb{P}^{D(N,m)}\to S\times\mathbb{P}^{D(N,m)}$
for some large $D(N,m)$. The subvariety $S\times(\mathbb{P}^{N})^{m}$ will sit inside $S\times\mathbb{P}^{D(N,m)}$ as a $\Phi_{N,m}$-invariant subvariety, so we will restrict our attention to $\Phi^{[m]}$ itself.
Let $\mathcal{X}_{m}$ denote the $m$-th fiber power of $\mathcal{X}$ over $S$, as a subvariety of $S\times(\mathbb{P}^{N})^{m}$, so each fiber of the projection $\mathcal{X}_{m}\to S$ is of the form $X_{s}^{m}$. Let $r_{m}^{N}$ be the relative special dimension of $\mathcal{X}_{m}$, so that
$r_{m}^{N}:=r_{\Phi^{[m]},\mathcal{X}_{m}}\leq mN=\dim\,(\mathbb{P}^{N})^{m}.$
If the set of $s\in S(\mathbb{C})$ such that $\mathcal{X}_{s}$ is $\Phi_{s}$-special is Zariski dense in $S$, then the $\Phi^{[m]}$-preperiodic points will be Zariski dense in $\mathcal{X}_{m}$ for all $m\geq 1$. Conjecture 1.2 implies that
$\dim\mathcal{X}_{m}\geq r_{m}^{N}$
for all $m\geq 1$. We will see, however, that $r_{m}^{N}=mN$ for all $m$, so that if $m>\dim S$, then
$r_{m}^{N}>\dim\mathcal{X}_{m}=\dim S+m\dim_{S}(\mathcal{X}),$
which is a contradiction. In other words, the conclusion of Conjecture 4.1 holds.
To see that $r_{m}^{N}=mN$, we work inductively on $m$ and appeal to a result of Fakhruddin [Fa2, Theorem 1.2]. He proved that any Zariski-closed subvariety $\mathcal{Y}\subset S\times\mathbb{P}^{N}$ which projects dominantly to $S$ and is invariant for $\Phi$ must either be all of $S\times\mathbb{P}^{N}$ or have codimension $N$. Note that there are no polarizable endomorphisms $\Psi$ commuting with $\Phi$ over all of $S$, except the iterates of $\Phi$ itself, so $\Phi$-special is the same as $\Phi$-preperiodic in this setting. Moreover, any Zariski-closed $\Phi$-invariant $\mathcal{Y}$ which projects dominantly to $S$ and has codimension $N$ must parameterize a finite collection of preperiodic points for $\Phi$; in particular, such a $\mathcal{Y}$ must project surjectively to $S$ and the fiber over every parameter in $S$ will have codimension $N$ in $\mathbb{P}^{N}$.
For $m=1$, we have $\mathcal{X}_{1}:=\mathcal{X}$ and by [Fa2, Theorem 1.2] the only $\Phi$-special subvariety containing $\mathcal{X}$ is $S\times\mathbb{P}^{N}$. So $r_{1}^{N}=N$. Now fix $m>1$, and assume that $r_{j}^{N}=jN$ for all $j<m$. Suppose that $\mathcal{Z}_{m}$ is a $\Phi^{[m]}$-special subvariety in $S\times(\mathbb{P}^{N})^{m}$ containing $\mathcal{X}_{m}$. We aim to show that $\mathcal{Z}_{m}=S\times(\mathbb{P}^{N})^{m}$.
Let
$p:S\times(\mathbb{P}^{N})^{m}\to S\times(\mathbb{P}^{N})^{m-1}$
denote the projection forgetting the last coordinate $\mathbb{P}^{N}$. By the induction hypothesis, we have $p(\mathcal{Z}_{m})=S\times(\mathbb{P}^{N})^{m-1}$, because $p(\mathcal{Z}_{m})$ must be $\Phi^{[m-1]}$-special and contain $\mathcal{X}_{m-1}$. Let $\mathcal{P}_{1}\subset S\times(\mathbb{P}^{N})^{m-1}$ be a subvariety of dimension $=\dim S$ which projects surjectively to $S$ and is pointwise fixed by $\Phi^{[m-1]}$. (In other words, $\mathcal{P}_{1}$ is a multisection of the projection from $S\times(\mathbb{P}^{N})^{m-1}$ to $S$, parameterizing a collection of fixed points.) Consider
$\mathcal{Z}^{x}_{m}:=p_{m}\left(p^{-1}(x)\cap\mathcal{Z}_{m}\right)\subset\mathbb{P}^{N},$
where $x$ is an element of $\mathcal{P}_{1}$, and $p_{m}:S\times(\mathbb{P}^{N})^{m}\to\mathbb{P}^{N}$ is the projection to the last coordinate. Note that
$\Phi^{[m]}(x,y)\;\in\;p^{-1}(x)\cap\Phi^{[m]}(\mathcal{Z}_{m})$
for every $x\in\mathcal{P}_{1}$ and $y\in\mathcal{Z}^{x}_{m}$, because $x$ is a fixed point. Since $\mathcal{Z}_{m}$ is preperiodic under $\Phi^{[m]}$, we see that $\mathcal{Z}^{x}_{m}$ defines a family of $\Phi$-preperiodic subvarieties in $\mathbb{P}^{N}$, as $x$ varies over $\mathcal{P}_{1}$. It follows from Fakhruddin’s theorem [Fa2, Theorem 1.2] that $\{\mathcal{Z}^{x}_{m}:x\in\mathcal{P}_{1}\}$ is either a family of points or all of $\mathbb{P}^{N}$ for all $x\in\mathcal{P}_{1}$.
In the former case, we deduce that $\operatorname{codim}\mathcal{Z}_{m}=N$ in $S\times(\mathbb{P}^{N})^{m}$, because the dimension of the fibers of $p|_{\mathcal{Z}_{m}}$ is upper semi-continuous over $S\times(\mathbb{P}^{N})^{m-1}$. It follows that $\mathcal{Z}_{m}$ defines a family of preperiodic points in $\mathbb{P}^{N}$ over all of $S\times(\mathbb{P}^{N})^{m-1}$ (viewed as an extended parameter space for the family of maps $\Phi_{s}$). But the fibers of the projection $p|_{\mathcal{Z}_{m}}$ over $\mathcal{X}_{m-1}$ must contain the corresponding fibers of $\mathcal{X}$, a contradiction.
We conclude that $\mathcal{Z}^{x}_{m}=\mathbb{P}^{N}$ for all $x\in\mathcal{P}_{1}$. Recall that we aim to show that $\mathcal{Z}_{m}=S\times(\mathbb{P}^{N})^{m}$. We now repeat this argument, replacing $\mathcal{P}_{1}$ with any family $\mathcal{P}$ of marked periodic points for $\Phi^{[m-1]}$ in $S\times(\mathbb{P}^{N})^{m-1}$ and work with an iterate of $\Phi$, and we deduce that the fibers of $\mathcal{Z}_{m}$ over $\mathcal{P}$ are also all of $\mathbb{P}^{N}$. As the periodic points of $\Phi^{[m-1]}$ are Zariski-dense in $S\times(\mathbb{P}^{N})^{m-1}$, and as the fibers of $p|_{\mathcal{Z}_{m}}$ are equal to $\mathbb{P}^{N}$ over this Zariski-dense set, we conclude that every fiber of $p|_{\mathcal{Z}_{m}}$ is equal to $\mathbb{P}^{N}$. Therefore, $\mathcal{Z}_{m}=S\times(\mathbb{P}^{N})^{m}$ and so the relative special dimension of $\mathcal{X}_{m}$ is $r_{m}^{N}=mN$. ∎
4.1. PCF density and sparsity in the moduli space
When $\mathcal{X}=\mathrm{Crit}(\Phi)$ is the critical locus of the family $\Phi$, then Conjecture 4.1 predicts the sparsity of postcritically finite maps in $\mathrm{M}_{d}^{N}$ for $N>1$, in the sense that the union of maps for which the critical hypersurface is preperiodic is contained in a proper Zariski-closed subset of the moduli space $\mathrm{M}_{d}^{N}$. This sparsity was conjectured in [IRS] and proved recently in [GTV].
For $N=1$, it is known that the postcritically finite maps form a Zariski-dense subset of the moduli space $\mathrm{M}_{d}^{1}$ of maps on $\mathbb{P}^{1}$, in any degree $d\geq 2$; see, for example, [BE, Main Theorem] or [De4, Theorem A]. This fact can also be viewed as a special case of Conjecture 1.1. To see this, suppose that $\Phi:S\times\mathbb{P}^{1}\to S\times\mathbb{P}^{1}$ is an algebraic family of endomorphisms of $\mathbb{P}^{1}$ degree $d>1$ for which the induced map $S\to\mathrm{M}_{d}^{1}$ is dominant. The cardinality of $\mathrm{Crit}(\Phi_{s})$ is $\leq 2d-2$ for every $s\in S$. This implies that
(4.1) $r^{1}_{m}:=r_{\Phi^{[m]},\mathrm{Crit}(\Phi)^{[m]}}\leq 2d-2$
for all $m\geq 1$, where $\mathrm{Crit}(\Phi)^{[m]}$ is the $m$-th fiber power of the critical locus, because the elements of the set $\mathrm{Crit}(\Phi_{s})^{m}\subset(\mathbb{P}^{1})^{m}$ have at most $22d-$ 2 distinct coordinate entries. (Note that the diagonal in $\mathbb{P}^{1}\times\mathbb{P}^{1}$ is invariant for the product map $(\Phi_{s},\Phi_{s})$.)
Recall that the bifurcation measure is nonzero on the moduli space $\mathrm{M}_{d}^{1}$; indeed, it was first proved in [BB, Proposition 6.3] that a rigid Lattès map must lie in the support of the measure. It follows that
(4.2) $\hat{T}_{\Phi^{[2d-2]}}^{\wedge(2d-2)}\wedge[\mathrm{Crit}(\Phi)^{[2d-2]}]\not=0,$
because this current projects to the bifurcation measure on $\mathrm{M}_{d}^{1}$ (via first projecting to the base $S$ and then via the natural map to $\mathrm{M}_{d}^{1}$); see [GTV, Proposition 1.4] for the computation relating the current in (4.2) to a bifurcation current in $S$. Let $\mathcal{X}$ be an irreducible component of $\mathrm{Crit}(\Phi)^{[2d-2]}$ for which $\hat{T}_{\Phi^{[2d-2]}}^{\wedge(2d-2)}\wedge[\mathcal{X}]\not=0$. Proposition 2.2 then implies that $r_{\Phi^{[2d-2]},\mathcal{X}}\geq 2d-2$ for this component; combined with (4.1), we have equality $r_{2d-2}^{1}=2d-2$. In particular, the coordinates of points in a fiber of $\mathcal{X}$ over $S$, in $(\mathbb{P}^{1})^{2d-2}$, are generally distinct. The implication (2) $\implies$ (1) of Conjecture 1.1 tells us that preperiodic points of $\Phi^{[2d-2]}$ are Zariski-dense in $\mathcal{X}$. But, over a Zariski-open and -dense subset $U$ of $S$ where the $22d-$ 2 critical points are distinct, the existence of a preperiodic point in $\mathcal{X}$ over $s\in U$ means that each of the $22d-$ 2 critical points for $\Phi_{s}$ is preperiodic. In other words, Conjecture 1.1 implies that the set of postcritically finite maps is Zariski-dense in $\mathrm{M}_{d}^{1}$.
5. Minimal $\Phi$-rank and stability
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of endomorphisms of degree $>1$. In this section, we introduce the notion of stability for subvarieties $\mathcal{X}\subset S\times\mathbb{P}^{N}$, following [GV]. We discuss the Geometric Dynamical Bogomolov Conjecture of Gauthier and Vigny [GV, Conjecture 1.9], which is a conjectural generalization of theorems proved in [GH1, CGHX] for abelian varieties over function fields of characteristic 0. It is related to our Conjecture 1.1 but is independent except in certain cases. We follow this discussion with a look at the special case of $J$-stability of $\Phi$ and the bifurcation current introduced in [BB] and studied further in [BBD].
5.1. Stability
Let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible subvariety of codimension $p$ which is flat over $S$. Gauthier and Vigny studied the currents $\hat{T}_{\Phi}^{\wedge k}\wedge[\mathcal{X}]$, for $k\geq 1$, in [GV] and defined an important notion: the subvariety $\mathcal{X}$ is said to be $\Phi$-stable if
$\hat{T}_{\Phi}^{N+1-p}\wedge[\mathcal{X}]=0;$
they prove that this is equivalent to the vanishing of the canonical ${\bf\Phi}$-height of ${\bf X}$ (for the height defined over the function field $\mathbb{C}(S)$, introduced by Gubler [Gu1, Gu2, Gu3]) [GV, Theorem B].
If $\mathcal{X}$ has codimension 1 in a $\Phi$-special variety $\mathcal{Y}$ but is not $\Phi$-special itself, so that $r_{\Phi,\mathcal{X}}=\dim_{S}\mathcal{X}+1$, then condition (2) of Conjecture 1.1 is exactly the condition of instability. The conjecture proposes that instability, in this case, is equivalent to the prevalence of $\Phi$-preperiodic points in $\mathcal{X}$.
But for general subvarieties $\mathcal{X}$, of arbitrary dimension, instability does not guarantee the existence of any preperiodic points, as the following example illustrates.
Example 5.1.
Let $S=\mathbb{C}$, and consider the family $\Phi:S\times\mathbb{P}^{2}\to S\times\mathbb{P}^{2}$ of degree $d=2$ defined by
$\Phi(s,x,y)=(s,x^{2}+s,y^{2}+s+10)$
in affine coordinates $(x,y)$ in $\mathbb{A}^{2}\subset\mathbb{P}^{2}$. Set
$\mathcal{X}=\{(s,0,0):s\in S\}.$
Then
$\hat{T}_{\Phi}=dd^{c}G\quad\mbox{ for }\quad G(s,x,y)=\max\{G_{s}(x),G_{s+10}(y)\}$
on $S\times\mathbb{C}^{2}$, where $G_{c}(z):=\lim_{n\to\infty}2^{-n}\log\max\{|f^{n}_{c}(z)|,1\}$ is the escape-rate function for the polynomial $f_{c}(z)=z^{2}+c$. It follows that
$\hat{T}_{\Phi}|_{\mathcal{X}}=dd^{c}G(s,0,0).$
In particular, since $s\mapsto G(s,0,0)$ is subharmonic, nonconstant, and bounded from below, we see that it cannot be harmonic, and therefore $\hat{T}_{\Phi}\wedge[\mathcal{X}]\not=0$. In other words, the subvariety $\mathcal{X}$ is unstable over $S$, in the sense of [GV]. On the other hand, we know that $G_{c}(0)=0$ if and only if $c$ is in the Mandelbrot set; it follows that $G(s,0,0)>0$ for all $s\in\mathbb{C}$, because there are no parameters $s$ where both $s$ and $s+10$ lie in the Mandelbrot set. In particular, there are no $\Phi$-preperiodic points in $\mathcal{X}$. Note that $\mathcal{X}$ is not contained in any nontrivial $\Phi$-special subvarieties of $S\times\mathbb{P}^{2}$, so that $r_{\Phi,\mathcal{X}}=2$. We can see immediately that
$\hat{T}_{\Phi}^{\wedge 2}\wedge[\mathcal{X}]=0$
because $\dim\mathcal{X}=1<2$, thus supporting the equivalence of Conjecture 1.1.
5.2. Minimal $\Phi$-rank
From Proposition 2.2, we see that a subvariety $\mathcal{X}\subset S\times\mathbb{P}^{N}$ is $\Phi$-stable if and only if it has minimal $\Phi$-rank, meaning that
$\operatorname{rank}_{\Phi}(\mathcal{X})=\dim_{S}\mathcal{X}.$
5.3. The Geometric Dynamical Bogomolov conjecture
Gauthier and Vigny formulated a conjecture that aims to characterize the stable subvarieties $\mathcal{X}\subset S\times\mathbb{P}^{N}$; that is, the subvarieties $\mathcal{X}$ of minimal $\Phi$-rank.
To formulate their conjecture in our terminology, we let $K=\mathbb{C}(S)$ be the function field of $S$ and introduce a few definitions. Recall that a family $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ is isotrivial if, after a base change $S^{\prime}\to S$, we can change coordinates by a family of automorphisms of $\mathbb{P}^{N}$ so that $\Phi_{s}$ becomes independent of the parameter $s\in S^{\prime}$. Similarly, a polarized endomorphism ${\bf\Phi}:{\bf Y}\to{\bf Y}$ defined over $\overline{K}$ is isotrivial if there exists a model family over $\mathbb{C}$ with an isotrivial extension to some $\mathbb{P}^{N}$. A subvariety ${\bf W}\subset{\bf Y}$ is $\mathbf{\Phi}$-isotrivial if it is itself independent of the parameter in the model over $\mathbb{C}$ after the coordinate change has been made. An irreducible subvariety $\mathcal{X}\subset S\times\mathbb{P}^{N}$ is said to come from an isotrivial factor of $\Phi$ if there exist integers $k_{1}\geq 1$ and $k_{2}\geq 0$, an isotrivial polarized endomorphism ${\bf\Psi}:{\bf Y}\to{\bf Y}$ with $\mathbf{\Psi}$-isotrivial subvariety ${\bf W}\subset{\bf Y}$, a subvariety $\mathbf{Z}\subset\mathbf{P^{N}}$ defined over $\overline{K}$ with $\mathbf{\Phi}^{k_{1}}(\mathbf{Z})=\mathbf{Z}$, and a dominant rational map $p:\mathbf{Z}\dashrightarrow\mathbf{Y}$ such that the following diagram commutes
${\mathbf{Z}}$${\mathbf{Z}}$${\mathbf{Y}}$${\mathbf{Y}}$$\scriptstyle{\mathbf{\Phi}^{k_{1}}}$$\scriptstyle{p}$$\scriptstyle{p}$$\scriptstyle{\mathbf{\Psi}}$
and so that $\mathbf{\Phi}^{k_{2}}(\mathbf{X})=p^{-1}(\mathbf{W})$.
Conjecture 5.2.
[GV] Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree $>1$, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible subvariety which is flat over $S$. The following are equivalent.
  1. (1)
    $\widehat{T}^{\wedge\dim_{S}(\mathcal{X})+1}_{\Phi}\wedge[\mathcal{X}]=0$
  2. (2)
    Either $\mathcal{X}$ is $\Phi$-special or it comes from an isotrivial factor.
As mentioned in §5.1, Gauthier and Vigny proved that condition (1) is equivalent to the vanishing of the canonical height $\hat{h}_{\bf\Phi}({\bf X})$ over the function field $K=\mathbb{C}(S)$ [GV, Theorem B]. Combining this with Gubler’s Inequality [Gu2, Lemma 4.1, Proposition 4.3], we know that condition (1) is also equivalent to the existence of a Zariski-dense set of geometrically small points in ${\bf X}$; that is, the existence of a generic sequence of points ${\bf x}_{n}\in{\bf X}(\overline{K})$ with $\hat{h}_{\bf\Phi}({\bf x}_{n})\to 0$ as $n\to\infty$.
That (2) implies (1) in Conjecture 5.2 is straightforward to prove, so the challenge is the conjectural (1) $\implies$ (2). This implication is known in dimension $N=1$ [De3, Theorem 1.1] and for subvarieties $\mathcal{X}$ with $\dim_{S}\mathcal{X}=0$ in arbitrary dimension $N$ [GV, Theorem A]; see also [CH1, CH2]. Gauthier–Vigny [GV, Theorem 29] proved that (1) implies (2) in Conjecture 5.2 when $\mathbf{\Phi}:\mathbf{P}^{2}\to\mathbf{P}^{2}$ is a non-isotrivial polynomial skew-product with an isotrivial first coordinate. Mavraki–Schmidt [MS, Theorem 4.1, Theorem 4.3], building on their work with Wilms [MSW], proved that Conjecture 5.2 holds when $\mathbf{\Phi}$ preserves a power of the projective line $(\mathbf{P}^{1})^{\ell}$ with the additional assumption that if $\ell\geq 3$, $\mathbf{\Phi}$ has no isotrivial factor. In the case of an abelian scheme, that is, where $\Phi$ preserves a family of abelian varieties $\mathcal{A}$ over $S$, inducing a homomorphism, and $\mathcal{X}$ is contained in $\mathcal{A}$, the conjecture was proved in [GH1, CGHX].
Note that Conjecture 5.2 and Conjecture 1.1 have distinct goals. Conjecture 5.2 is a characterization of subvarieties $\mathcal{X}$ of minimal $\Phi$-rank, i.e., $\operatorname{rank}_{\Phi}(\mathcal{X})=\dim_{S}\mathcal{X}$, while Conjecture 1.1 is a characterization of subvarieties $\mathcal{X}$ of maximal $\Phi$-rank, i.e., $\operatorname{rank}_{\Phi}(\mathcal{X})=r_{\Phi,\mathcal{X}}$. (Recall that we gave bounds on $\operatorname{rank}_{\Phi}(\mathcal{X})$ in Proposition 2.2.) A subvariety $\mathcal{X}$ will have both maximal and minimal $\Phi$-rank if and only if $\dim_{S}\mathcal{X}=r_{\Phi,\mathcal{X}}$, meaning that $\mathcal{X}$ is itself $\Phi$-special. And in that case, the subvariety $\mathcal{X}$ contains a Zariski-dense set of persistently preperiodic points over $S$, giving rise to a Zariski-dense set of geometric preperiodic points in $\mathbf{X}$.
There is interesting overlap of the two conjectures in the case where $\dim_{S}\mathcal{X}=r_{\Phi,\mathcal{X}}-1$; that is, where $\mathcal{X}$ is a family of hypersurfaces within a $\Phi$-special subvariety. In this case, Conjecture 5.2 combined with the Dynamical Manin-Mumford conjecture of §3.1 (the $\dim S=0$ case of Conjecture 1.1) implies the unknown implication (1) $\implies$ (2) of Conjecture 1.1 for arbitrary $S$; the DMM conjecture is used to eliminate the possibility of an isotrivial factor which is not itself $\Phi$-special but which has a Zariski-dense set of preperiodic points. But Conjecture 1.1 does not seem to imply the unknown implication (1) $\implies$ (2) of Conjecture 5.2, because it does not classify subvarieties where Zariski-density of preperiodic points fails. See Proposition 5.3 below for a special case.
5.4. $J$-Stability in any dimension
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree $>1$ and $\mathrm{Crit}(\Phi)$ the critical locus in $S\times\mathbb{P}^{N}$. In this case, condition (2) of Conjecture 1.1 can be interpreted in terms of $J$-stability of the family, as mentioned in §3.2 in the case $N=1$. Note that the relative dimension of $\mathrm{Crit}(\Phi)$ over $S$ is $N-1$, so that the relative special dimension $r_{\Phi,\mathcal{X}}$ is either $N-1$ or $N$ for each irreducible component $\mathcal{X}$ of $\mathrm{Crit}(\Phi)$.
By definition, the family $\Phi$ is $J$-stable, as defined by Berteloot, Bianchi, and Dupont in [BBD], if
$\hat{T}_{\Phi}^{\wedge N}\wedge[\mathrm{Crit}(\Phi)]=0.$
The bifurcation current is equal to the push-forward of $\hat{T}_{\Phi}^{\wedge N}\wedge[\mathrm{Crit}(\phi)]$ to the base $S$; see [BB]. So if $r_{\Phi,\mathcal{X}}=N$ for a component $\mathcal{X}$ of the critical locus, the nonvanishing of the current $\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}\wedge[\mathcal{X}]$ in condition (2) of Conjecture 1.1 implies instability of the family.
If $r_{\Phi,\mathcal{X}}=N-1$ for some component $\mathcal{X}$ of $\mathrm{Crit}(\Phi)$, the associated family of critical hypersurfaces will be $\Phi$-special, and so contain a Zariski-dense set of preperiodic points. Moreover, the current $\hat{T}_{\Phi}^{N-1}\wedge[\mathcal{X}]$ is non-vanishing for dimension reasons, by Lemma 2.3. So the equivalence of (1) and (2) in Conjecture 1.1 holds for this component.
In [BBD], the authors proved that Misiurewicz parameters are dense in the support of the bifurcation current. A Misiurewicz parameter $s_{0}\in S$ is one where a component of the critical locus of $\Phi_{s_{0}}$ contains a repelling periodic point in the support of its measure of maximal entropy so that the graph of this point over $S$ (defined in a neighborhood of $s_{0}$) has a proper intersection with $\mathrm{Crit}(\Phi)$; see [BBD, Definition 1.5]. Working through the details of their proof, we see that they in fact prove that preperiodic points are dense in the support of $\hat{T}_{\Phi}^{\wedge N}\wedge[\mathrm{Crit}(\Phi)]$, and so Zariski-dense in $\mathrm{Crit}(\Phi)$ itself, thus proving the implication (2) $\implies$ (1) of Conjecture 1.1 in this setting. More precisely, for any parameter $s_{0}\in S$ in the support of the bifurcation current, they show in the proof of [BBD, Theorem 1.6 and Proposition 3.8] that the forward orbit of the critical locus will intersect the graphs of a collection of repelling cycles moving holomorphically near $s_{0}$. The motion comes from a hyperbolic set in the Julia set over $s_{0}$, and this set of graphs is Zariski dense in $S\times\mathbb{P}^{N}$. If the parameter $s_{0}$ is in the projection of the support of $\hat{T}_{\Phi}^{\wedge N}\wedge[\mathcal{X}]$ for a component $\mathcal{X}$, then it is an iterate of $\mathcal{X}$ which intersects a positive-measure collection of leaves of this motion (for any appropriate measure on the hyperbolic set). It follows that the preperiodic points are Zariski-dense in $\mathcal{X}$ itself. The proof of Theorem 1.3 that we provide in Section 6 follows this same strategy.
But the converse implication of Conjecture 1.1, that (1) $\implies$ (2) for all components $\mathcal{X}$ of $\mathrm{Crit}(\Phi)$, remains open for every $N>1$.
Proposition 5.3.
Conjecture 1.1 implies that an algebraic family $\Phi$ of maps on $\mathbb{P}^{N}$ of degree $d>1$ is $J$-stable if and only if each irreducible component $\mathcal{X}$ of the critical locus $\mathrm{Crit}(\Phi)$ is either $\Phi$-special or the set $\mathrm{Prep}(\Phi)\cap\mathcal{X}$ is contained in a proper algebraic subvariety of $\mathcal{X}$.
To be clear: we know that if $\Phi$ is not $J$-stable, and if $\mathcal{X}$ is a component of $\mathrm{Crit}(\Phi)$ for which $\hat{T}_{\Phi}^{\wedge N}\wedge[\mathcal{X}]\not=0$, then the $\Phi$-preperiodic points are Zariski dense in $\mathcal{X}$. This follows from the work of [BBD]. But we do not know that Zariski-density of (non-persistent) preperiodic points in a component $\mathcal{X}$ of $\mathrm{Crit}(\Phi)$ is enough to guarantee instability, except in dimension $N=1$ (see §3.2), because we do not know a priori that the points are repelling.
Moreover, for $N>1$, we do not have a classification of families $\Phi$ for which the latter condition of Proposition 5.3 happens, i.e., when $\mathrm{Prep}(\Phi)\cap\mathcal{X}$ is not Zariski-dense in a component $\mathcal{X}$ of $\mathrm{Crit}(\Phi)$. The Geometric Bogomolov Conjecture 5.2 above aims to rectify this, proposing that such an $\mathcal{X}$ must come from an isotrivial factor.
6. Forced intersections: the proof of Theorem 1.3
Throughout this section, we will assume that condition (2) of Conjecture 1.1 holds. That is, we let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree $>1$ and $\mathcal{X}\subset S\times\mathbb{P}^{N}$ an irreducible subvariety which is flat over $S$. Let $\mathcal{Y}$ be a $\Phi$-special subvariety for $\Phi$ of relative dimension $r_{\Phi,\mathcal{X}}$ that contains $\mathcal{X}$, which is preperiodic for an endomorphism $\Psi$ that commutes with $\Phi$ on $\mathcal{Y}$. We assume that
$T:=\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}\wedge[\mathcal{X}]\not=0$
as a current on $S\times\mathbb{P}^{N}$. Recall that the slices of $\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}$ over $S$ are measures of maximal entropy for the restriction of $\Psi$ to (an iterate of) $\mathcal{Y}$. Recall also that $\Phi$ and $\Psi$ have the same sets of preperiodic points in $\mathcal{Y}$. Working with an iterate of $\Psi$ instead of $\Phi$, and with forward iterates of $\mathcal{X}$ and $\mathcal{Y}$, we will assume that $\mathcal{Y}$ is fixed by $\Phi$ itself, so that each vertical slice of $\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}$ is the measure of maximal for $\Phi$ in $\mathcal{Y}_{s}$, for $s\in S$.
To prove that (2) $\implies$ (1) in Conjecture 1.1, we follow the proof of [Du, Theorem 0.1], which was extended to maps on $\mathbb{P}^{N}$ in [Ga1, Theorem 2.2].
Consider the nonzero current $T_{S}=\pi_{*}T$ on $S$, where $\pi:S\times\mathbb{P}^{N}\to S$ is the projection. Fix $\lambda_{0}$ in the support of $T_{S}$, and choose a hyperbolic set $K_{0}$ for $\Phi_{\lambda_{0}}$ in $\mathcal{Y}_{\lambda_{0}}$ which moves holomorphically over a neighborhood $U$ of $\lambda_{0}$ in $S$; we select $K_{0}$ so that it supports a probability measure $\nu$ with maximal entropy for the restriction of $\Phi_{\lambda_{0}}$ to $K_{0}$ and which is a measure of type PLB. See, for example, [DS1, Theorems 3.2.1 and 3.9.5], [BBD, Theorems 3.9 and 3.11], and [Ga1, Proposition 1.5]. In particular, $\nu$ puts no mass on proper analytic subvarieties of $\mathcal{Y}_{\lambda_{0}}$, and so its support is Zariski dense in $\mathcal{Y}_{\lambda_{0}}$. Let
$r:=r_{\Phi,\mathcal{X}}=\dim_{S}\mathcal{Y}$
denote the relative dimension of $\mathcal{Y}$. For each $z\in K_{0}$, let $\Gamma_{z}$ denote the graph of the motion of $z$ over the open set $U$ in $S$, and define
[…]
Then $\hat{\nu}$ is a positive and uniformly laminar $(r,r)$-current on
$\mathcal{Y}_{U}:=\pi^{-1}(U)\cap\mathcal{Y}.$
Observe that
$\frac{1}{d^{nr}}(\Phi|_{\mathcal{Y}}^{n})^{*}\hat{\nu}\longrightarrow\hat{T}_{\Phi}^{\wedge r}$
by the characterization of $\hat{T}_{\Phi}$ and the continuity of its potentials. On the other hand, we know that $\hat{\nu}=dd^{c}V$ for a locally bounded current of type $(r-1,r-1)$ since $\nu$ is a PLB measure, as explained in the proof of [Ga1, Theorem 2.6]. It follows that
$\frac{1}{d^{nr}}(\Phi^{n})^{*}\hat{\nu}\wedge[\mathcal{X}]\longrightarrow\hat{T}_{\Phi}^{\wedge r}\wedge[\mathcal{X}]=T>0.$
Now consider the iterated images of $\mathcal{X}$ over $U$; let $\mathcal{X}_{n}=\Phi^{n}(\mathcal{X})$. Since $T\not=0$, we know that $\dim\mathcal{X}\geq r$; let $\chi_{U}$ be a smooth function with compact support in $U$ which is $\equiv 1$ in a neighborhood of $\lambda_{0}$. Set
$\beta=(\chi_{U}\circ\pi)\,\hat{\omega}^{\dim\mathcal{X}-r}$
on $\mathcal{Y}_{U}$, where $\hat{\omega}$ is a smooth $(1,1)$-form on $S\times\mathbb{P}^{N}$ restricting to Fubini-Study on each fiber $\mathbb{P}^{N}$. Then
$\int_{\mathcal{Y}_{U}}(\Phi^{n})^{*}\hat{\nu}\wedge[\mathcal{X}]\wedge\beta$ $=$ $\int_{\mathcal{Y}_{U}}\hat{\nu}\wedge(\Phi^{n})_{*}[\mathcal{X}]\wedge(\Phi^{n})_{*}\beta$
$=$ $c_{n}\,\int_{\mathcal{Y}_{U}}\hat{\nu}\wedge[\mathcal{X}_{n}]\wedge(\Phi^{n})_{*}\beta$
for some $c_{n}>0$. In particular, the current $\hat{\nu}\wedge[\mathcal{X}_{n}]$ will be nonzero in $\mathcal{Y}_{U}$ for all sufficiently large $n$. But by [Du, Theorem 3.1], we know that
$\hat{\nu}\wedge[\mathcal{X}_{n}]=\int_{K_{0}}[\Gamma_{z}]\wedge[\mathcal{X}_{n}]\,d\nu,$
and the intersections of $\mathcal{X}_{n}$ with a positive $\nu$-measure set of graphs $\Gamma_{z}$ will be transverse, over a small neighborhood of $\lambda_{0}$. Since the repelling periodic points are dense in $K_{0}$, it follows that $\mathcal{X}_{n}$ must intersect the graphs of repelling points, in a set which is dense in a set of positive $\nu$-measure. Moreover, these intersections are Zariski dense in $\mathcal{X}_{n}$. We conclude that the preperiodic points are Zariski dense in $\mathcal{X}$.
This completes the proof of Theorem 1.3. ∎